After it had cooled the dry marble and test tube weighed 17.56 g. I know it is difficult to see the tick marks for the volume of gas in the video so by the end I collected 35.3 mL of CO2. Calculate the theoretical yield of CO2 in liters (do a stoichiometry problem use three decimals)


1
Expert's answer
2021-03-05T03:27:12-0500

CaCO3 + 2HCl = H2O + CO2 + CaCl2

n (CO2) = n (CaCO3)

n (CO2) = n (CaCO3) = V/22.4 = 0.0353/22.4 = 0.002 mol